Solve the system: 2x + 3y = 18
4x – 5y = 14

Answer by checkley77(12844)

2x+3y=18 multiply by -2 & add
4x–5y=14
-4x-6y=-36
-------------
-11y=-22
y=-22/-11
y=2 ans.
2x+3*2=18
2x+6=18
2x=18-6
2x=12
x=12/2
x=6 ans.